AASTEX, AMS math, and L^{A}TEX symbols Table 1: Additional AASTEX symbols

≲ \lesssim, \la ≳  \gtrsim, \ga
μm \micron                \sbond
\dbond                 \tbond
⊙ \sun  ⊕  \earth
/ ◯ \diameter
∘  \arcdeg, \degr ⊓ ⊔  \sq
′  \arcmin         ′′  \arcsec
.^{d}  \fd  .^{h}  \fh
.^{m} \fm  .^{s}  \fs
.^{∘}  \fdg  .′  \farcm
.′′  \farcs  .^{p}  \fp
\frac{1}{2} \onehalf  UBVR \ubvr
\frac{1}{3} \onethird  U −B \ub
\frac{2}{3} \twothirds  B−V \bv
\frac{1}{4} \onequarter  V −R \vr
\frac{3}{4} \threequarters U −R \ur

Table 2: Text-mode accents

ò \`{o}  ō \={o}  oo \t{oo}
ó \'{o}  ȯ \.{o}  o̧ \c{o}
ô \^{o}  ŏ \u{o}        o. \d{o}
ö \"{o}  ǒ \v{o}        ō \b{o}
õ \~{o}  ő \H{o}

Table 3: National symbols

œ \oe  å \aa  ł \l
Œ \OE  Å \AA  Ł \L
æ \ae  ø \o  ß \ss
Æ \AE  Ø \O

Table 4: Math-mode accents

â \hat{a}          ȧ \dot{a}
ǎ \check{a}        ä \ddot{a}
ã \tilde{a}        ă \breve{a}
á \acute{a}        ā \bar{a}
à \grave{a}       a⃗ \vec{a}

Table 5: Greek and Hebrew letters (math mode)

   α \alpha  ν \nu
   β \beta  ξ \xi
   γ \gamma  o o
   δ \delta  π \pi
   ϵ \epsilon  ρ \rho
   ζ \zeta  σ \sigma
   η \eta  τ \tau
   θ \theta  υ \upsilon
   ι \iota  ϕ \phi
   κ \kappa  χ \chi
   λ \lambda  ψ \psi
   μ \mu  ω \omega
   𝟋 \digamma  κ \varkappa
   ε \varepsilon  ς \varsigma
   ϑ \vartheta  φ \varphi
   ϱ \varrho
   Γ \Gamma  Σ \Sigma
   Δ \Delta  Υ \Upsilon
   Θ \Theta  Φ \Phi
   Λ \Lambda  Ψ \Psi
   Ξ \Xi  Ω \Omega
   Π \Pi
   ℵ \aleph  ℶ \beth
   ℷ \gimel  ℸ \daleth

Table 6: Binary operators (math mode)

± \pm  ∩ \cap
∓ \mp  ∪ \cup
\ \setminus  ⊎ \uplus
· \cdot  ⊓ \sqcap
× \times  ⊔ \sqcup
∗ \ast  ◁ \triangleleft
⋆ \star  ▷ \triangleright
⋄ \diamond  ≀ \wr
∘ \circ  ◯ \bigcirc
• \bullet  △ \bigtriangleup
÷ \div  ▽ \bigtriangledown
⊲ \lhd  ⊳ \rhd
∨ \vee  ⊙ \odot
∧ \wedge  † \dagger
⊕ \oplus  ‡ \ddagger
⊖ \ominus  ⨿ \amalg
⊗ \otimes  ⊴ \unlhd
⊘ \oslash  ⊵ \unrhd

A1 AASTEX, AMS math, and L^{A}TEX symbols Table 7: AMS binary operators (math mode)

∔ \dotplus  ⋉ \ltimes
∖ \smallsetminus  ⋊ \rtimes
⋒ \Cap, \doublecap  ⋋ \leftthreetimes
⋓ \Cup, \doublecup  ⋌ \rightthreetimes
⊼ \barwedge  ⋏ \curlywedge
⊻ \veebar  ⋎ \curlyvee
⩞ \doublebarwedge
⊟ \boxminus  ⊝ \circleddash
⊠ \boxtimes  ⊛ \circledast
⊡ \boxdot  ⊚ \circledcirc
⊞ \boxplus  ▪ \centerdot
⋇ \divideontimes  ⊺ \intercal

Table 8: Miscellaneous symbols

† \dag  § \S
\begin{matrix}c \\ ◯ \end{matrix} \copyright  ‡ \ddag
¶ \P  £ \pounds
# \#  $ \$
% \%  & \&
\_{ \{
} \}

Table 9: Miscellaneous symbols (math mode)

ℵ \aleph  ′ \prime
ℏ \hbar  ∅ \emptyset
ı \imath  ∇ \nabla
ȷ \jmath  √ \surd
ℓ \ell  ⊤ \top
℘ \wp  ⊥ \bot
ℜ \Re  ∥ \|
ℑ \Im  ∠ \angle
∂ \partial  △ \triangle
∞ \infty  \ \backslash
□ \Box  ◇ \Diamond
∀ \forall  ♯ \sharp
∃ \exists  ♣ \clubsuit
¬ \neg  ♢ \diamondsuit
♭ \flat  ♡ \heartsuit
♮ \natural  ♠ \spadesuit
℧ \mho

Table 10: AMS miscellaneous symbols (math mode)

ℏ \hbar  ‵ \backprime
ℏ \hslash  ∅ \varnothing
△ \vartriangle  ▲ \blacktriangle
▽ \triangledown  ▼ \blacktriangledown
□ \square  ■ \blacksquare
♢ \lozenge  ♦ \blacklozenge
Ⓢ \circledS  ⋆ \bigstar
∠ \angle  ∢ \sphericalangle
∡ \measuredangle
∄ \nexists  ∁ \complement
℧ \mho  ð \eth
Ⅎ \Finv  ⧸ \diagup
⅁ \Game  ⧹ \diagdown
𝕜 \Bbbk  ↾ \restriction

Table 11: Arrows (math mode)

← \leftarrow  ⟵ \longleftarrow
⇐ \Leftarrow  ⟸ \Longleftarrow
→ \rightarrow  ⟶ \longrightarrow
⇒ \Rightarrow  ⟹ \Longrightarrow
↔ \leftrightarrow  ⟷ \longleftrightarrow
⇔ \Leftrightarrow  ⟺ \Longleftrightarrow
↦ \mapsto  ⟼ \longmapsto
↩ \hookleftarrow  ↪ \hookrightarrow
↼ \leftharpoonup  ⇀ \rightharpoonup
↽ \leftharpoondown  ⇁ \rightharpoondown
⇌ \rightleftharpoons ⇝ \leadsto
↑ \uparrow  ⇕ \Updownarrow
⇑ \Uparrow  ↗ \nearrow
↓ \downarrow  ↘ \searrow
⇓ \Downarrow  ↙ \swarrow
↕ \updownarrow  ↖ \nwarrow

A2 AASTEX, AMS math, and L^{A}TEX symbols Table 12: AMS arrows (math mode)

←−− \dashleftarrow  −−→ \dashrightarrow
⇇ \leftleftarrows  ⇉ \rightrightarrows
⇆ \leftrightarrows  ⇄ \rightleftarrows
⇚ \Lleftarrow  ⇛ \Rrightarrow
↞ \twoheadleftarrow  ↠ \twoheadrightarrow
↢ \leftarrowtail  ↣ \rightarrowtail
↫ \looparrowleft  ↬ \looparrowright
⇋ \leftrightharpoons  ⇌ \rightleftharpoons
↶ \curvearrowleft  ↷ \curvearrowright
⟲ \circlearrowleft  ⟳ \circlearrowright
↰  \Lsh  ↱  \Rsh
⇈ \upuparrows  ⇊ \downdownarrows
↿  \upharpoonleft  ↾  \upharpoonright
⇃  \downharpoonleft  ⇂  \downharpoonright
⊸ \multimap  ⇝ \rightsquigarrow
↭ \leftrightsquigarrow
↚ \nleftarrow  ↛ \nrightarrow
⇍ \nLeftarrow  ⇏ \nRightarrow
↮ \nleftrightarrow  ⇎ \nLeftrightarrow

Table 13: Relations (math mode)

≤ \leq  ≥ \geq
≺ \prec  ≻ \succ
⪯ \preceq  ⪰ \succeq
≪ \ll  ≫ \gg
⊂ \subset  ⊃ \supset
⊆ \subseteq  ⊇ \supseteq
⊏ \sqsubset  ⊐ \sqsupset
⊑ \sqsubseteq  ⊒ \sqsupseteq
∈ \in  ∋ \ni
⊢ \vdash  ⊣ \dashv
⌣ \smile  | \mid
⌢ \frown  ∥ \parallel
≠ \neq  ⊥ \perp
≡ \equiv             ≅ \cong
∼ \sim  ▷◁ \bowtie
≃ \simeq  ∝ \propto
≍ \asymp  ⊧ \models
≈ \approx            ≐ \doteq
⋈ \Join

Table 14: AMS binary relations (math mode)

≦ \leqq  ≧ \geqq
⩽ \leqslant  ⩾ \geqslant
⪕ \eqslantless  ⪖ \eqslantgtr
≲ \lesssim  ≳ \gtrsim
⪅ \lessapprox  ⪆ \gtrapprox
≊ \approxeq  ≂ \eqsim
⋖ \lessdot  ⋗ \gtrdot
≪ \lll, \llless  ≫ \ggg, \gggtr
≶ \lessgtr  ≷ \gtrless
⋚ \lesseqgtr  ⋛ \gtreqless
⪋ \lesseqqgtr  ⪌ \gtreqqless
≑ \doteqdot, \Doteq  ≖ \eqcirc
≓ \risingdotseq  ≗ \circeq
≒ \fallingdotseq  ≜ \triangleq
∽ \backsim  ∼ \thicksim
⋍ \backsimeq  ≈ \thickapprox
⫅ \subseteqq  ⫆ \supseteqq
⋐ \Subset  ⋑ \Supset
⊏ \sqsubset  ⊐ \sqsupset
≼ \preccurlyeq  ≽ \succcurlyeq
⋞ \curlyeqprec  ⋟ \curlyeqsucc
≾ \precsim  ≿ \succsim
≾ \precapprox  ≿ \succapprox
◁ \vartriangleleft  ▷ \vartriangleright
⊴ \trianglelefteq  ⊵ \trianglerighteq
⊨ \vDash  ⊩ \Vdash
⊪ \Vvdash
⌣ \smallsmile  ∣  \shortmid
⌢ \smallfrown  ∥  \shortparallel
≏ \bumpeq  ≬  \between
≎ \Bumpeq  ⋔ \pitchfork
∝ \varpropto  ϶  \backepsilon
◀ \blacktriangleleft  ▶ \blacktriangleright
∴ \therefore  ∵ \because

A3 AASTEX, AMS math, and L^{A}TEX symbols

Table 15: AMS negated relations (math mode)

≮ \nless  ≯ \ngtr
≰ \nleq  ≱ \ngeq
⩽̸ \nleqslant  ⩾̸ \ngeqslant
≦̸ \nleqq  ≧̸ \ngeqq
⪇ \lneq  ⪈ \gneq
≨ \lneqq  ≩ \gneqq
≨ \lvertneqq  ≩ \gvertneqq
⋦ \lnsim  ⋧ \gnsim
⪉ \lnapprox  ⪊ \gnapprox
⊀ \nprec  ⊁ \nsucc
⪯̸ \npreceq  ⪰̸ \nsucceq
⪵ \precneqq  ⪶ \succneqq
⋨ \precnsim  ⋩ \succnsim
⪹ \precnapprox  ⪺ \succnapprox
≁ \nsim  ≇ \ncong
∤ \nshortmid  ∦ \nshortparallel
∤ \nmid  ∦ \nparallel
⊬ \nvdash  ⊭ \nvDash
⊮ \nVdash  ⊯ \nVDash
⋪ \ntriangleleft  ⋫ \ntriangleright
⋬ \ntrianglelefteq  ⋭ \ntrianglerighteq
⊈ \nsubseteq  ⊉ \nsupseteq
⫅̸ \nsubseteqq  ⫆̸ \nsupseteqq
⊊ \subsetneq  ⊋ \supsetneq
⊊ \varsubsetneq  ⊋ \varsupsetneq
⫋ \subsetneqq  ⫌ \supsetneqq
⫋ \varsubsetneqq  ⫌ \varsupsetneqq

Table 16: Variable-sized symbols (math mode)

  ∑ ∑ \sum  ⋂ ⋂ \bigcap
  ∏ ∏ \prod  ⋃ ⋃ \bigcup
  ∐ ∐ \coprod  ⨆ ⨆ \bigsqcup
  ∫ ∫ \int  ⋁ ⋁ \bigvee
  ∮ ∮ \oint  ⋀ ⋀ \bigwedge
  ⨀ ⨀ \bigodot  ⨂ ⨂ \bigotimes
  ⨁ ⨁ \bigoplus  ⨄ ⨄ \biguplus

Table 17: Delimiters (math mode)

                                                           (  (  )  )
                                                           [  [  ]  ]
                                                           {  \{  }  \}
                                                           ⌊  \lfloor  ⌋  \rfloor
                                                           ⌈  \lceil  ⌉  \rceil
                                                           ⟨  \langle  ⟩  \rangle
                                                           /  /  \ \backslash
                                                           |  \vert  ∥  \Vert
                                                           ↑  \uparrow  ⇑  \Uparrow
                                                           ↓ \downarrow  ⇓ \Downarrow
                                                           ↕ \updownarrow  ⇕ \Updownarrow
                                                           ⌜ \ulcorner  ⌝  \urcorner
                                                           ⌞ \llcorner  ⌟  \lrcorner

Table 18: Function names (math mode)

                                                            \arccos   \csc    \ker      \min
                                                            \arcsin   \deg    \lg       \Pr
                                                            \arctan   \det    \lim      \sec
                                                            \arg      \dim    \liminf   \sin
                                                            \cos      \exp    \limsup   \sinh
                                                            \cosh     \gcd    \ln       \sup
                                                            \cot      \hom    \log      \tan
                                                            \coth     \inf    \max      \tanh

A4